OpenStudy (dmezzullo):

Rose worked the following problem incorrectly. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1 30 ÷ 5(3 − 1) − 4 Line 2 30 ÷ 5(2) − 4 Line 3 30 ÷ 10 − 4 Line 4 3 − 4 Line 5 −1

OpenStudy (dmezzullo):

neeed help plzzzzzz

OpenStudy (dmezzullo):

can anyone help me?!?!?!?!

OpenStudy (dmezzullo):

can someone plase help me??????????

OpenStudy (dmezzullo):

anyone?

OpenStudy (anonymous):

Rose takes a very very correct approach. But you don't. I told you before.

OpenStudy (andriod09):

It's line 3, the correct answer would be \[30\div5(2)-4\]\[30\div5\] then to do: \[5(2)\]

OpenStudy (cwrw238):

i'm not sure of this one - wolfram gives the answer = 8 it does the division 30/5 first = 6 then 6*2 - 4 = 8

OpenStudy (cwrw238):

i'm confused!!!

OpenStudy (cwrw238):

@hartnn - whats your opinion on this?

OpenStudy (turingtest):

my opinion is that this is why god created the fractional division symbol :P

OpenStudy (andriod09):

oki, time for me to step in here. The correct way for this problem to be solved is this: Step 1, parenthesis \[30 ÷ 5(3 − 1) − 4\]\[30\div5(2)-4\]Step 2, Exponets, but there aren't any, so lets skip that step Step 3, Multiplication/division \[30 ÷ 5(3 − 1) − 4\]\[6(2)-4\]\[12-4\] Step 4, addition/subtraction: \[21-4\] \[8\] to remember these steps try this: Parenthesis Please Exponets Excuse Multiplication My Division Dear Addition Aunt Subtraction Sally

hartnn (hartnn):

it should be 8, 30÷5(2)-4 is same as 30÷5*2 -4 while evaluating such expressions, where there is equal priority to ÷ and * , we always go from left to right. so 12-4=8

OpenStudy (cwrw238):

aahh yes

OpenStudy (cwrw238):

what confused me was the 5(2) - i was considering that to be the parentheses part